2.2 in fraction form

Found 2 solutions by Vladdroid, fcabanski:
Answer by Vladdroid(91)
Answer by fcabanski(1391)

2 and 2 tenths is 2 + 2/10 = 2 + 1/5 = 2 and 1/5. Convert into an improper fraction by multiplying the denominator by the whole number, adding the numerator, and placing that result over the denominator.


5x2+1 = 10+1 = 11/5.

Hope the solution helped. Sometimes you need more than a solution. Contact fcabanski@hotmail.com for online, private tutoring, or personalized problem solving (quick for groups of problems.)